A drop of water of radius 6 mm breaks down into number of droplets each of radius 1mm how many droplets will be formed

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Volume of large water drop
= (4/3)*pi* 6³
Now
This should be equal to n drops of 1 mm radius
Hence
(4/3)*pi* 6³ = n* (4/3)* pi* 1³
n = 216
 
Hence 216 drops will be formed.
 
Regards
Arun (askIITians forum expert)